OpenStudy (anonymous):

PLEASE HELP!!!! Solve for x. square root of x plus twenty-one + square root of x = 7

OpenStudy (priyar):

u mean..\[\sqrt{x+21} +\sqrt{x} = 7\]

OpenStudy (priyar):

take \[\sqrt{x}\] to thje other side and square on both sides

OpenStudy (priyar):

*the

OpenStudy (priyar):

@kawaiidayna ?

OpenStudy (anonymous):

Nevermind! I got it!!